The disk-shaped pulley of mass 10.0kg has a radius R and a moment of inertia I turns on a frictionless bearing. The rope does not slip over the pulley.

a) Find the work done by the frictional force, if the system drops 0.60m the coefficient of kinetic friction between the table and the 5.00kg block is 0.35

b) How fast is the system moving after dropping 0.6m?

5.00kg 10Kg 9.00kg
Answer #1

Free body diagram Shout 59,= T, -0.35859 sa,=T-4R, R=593 TUR Tisg a,=a2 922=gg-T2 1. . I q = TER-TIXR Cookg, a » į mR 2 X 9 =